# Trade-off of Security, Latency, and Throughput of the Nakamoto Consensus

### v3

@@ 4.2. Throughput and Related Trade-offs
Empirically, the block propagation delay is approximately an affine function of the block size (decker2013information, ; fechner2022calibrating, ) (even with a coded design (zhang2022speeding, )). That is, the propagation delay bound can be modeled as
(21) $\Delta(B)=\frac{B}{r}+\nu$
where $B$ is the block size in KB and $r$ and $\nu$ are constants. Here, $r$ represents a fundamental limitation on the end-to-end transmission rate of the peer-to-peer network. Based on (20), this rate and the desired fault tolerance determines an upper bound for the optimistic throughput:
(22) $\frac{\lambda B}{1+\lambda\Delta}<\frac{1-2\beta}{1-\beta-\beta^{2}}\,r.$
With a close upper bound on the probability of safety violation as a function of the system parameters and the confirmation latency, e.g., in the form of $p(\lambda\Delta,\beta,k)$ where a user may select the confirmation depth $k$ to achieve a desired safety. It is well understood that a user making a very large transaction is generally willing to tolerate a longer confirmation latency. The theorems in this paper allow this choice to be made much more accurately than before.
@@
(23a) $\text{maximize}_{\lambda,B,k}\quad$ $\frac{\lambda B}{1+\lambda\,\Delta(B)}$
(23b) subject to $p(\lambda\,\Delta(B),\beta,k)\leq q$
(23c) $k\leq\frac{(1-\beta)\lambda}{1+(1-\beta)\lambda\Delta}d$
where $q$ denotes the maximum tolerable probability of violation and $d$ denotes the expected confirmation time. Here (23c) constrains $k$ such that at the slowest growth rate, gaining $k$ heights takes no more than $d$ units of time on average. In Sec. 6, we give such a numerical example under assumption (21) with empirically determined constants.
4.3. Bounds for Given Confirmation Time
Theorem 8 (Upper bound for given confirmation time).
Suppose the mining rates and the delay upper bound satisfy (1). Given a sufficiently large $s$, the safety of a transaction which becomes public during $(s,s+2\Delta]$ and is confirmed by time $t$ can be violated with a probability no higher than
(24) $\inf_{n\in\{1,2,\dots\}}\{1-F(n-1)+G(n,t)\},$
@@ 4.3. Bounds for Given Confirmation Time
(25) $\begin{split}&G(n,t)=\sum_{i=0}^{\infty}\sum_{j=0}^{\infty}\sum_{k=0}^{\infty}f_{1}(i;a(t+2\Delta))e(j)e(k)\\ &\times(1-F_{2}(t-(i+j+k+n+2)\Delta;i+j+k+n,h))\end{split}$
@@
(26) $\begin{split}\overline{G}(n,t)&=1-F_{1}(i_{0};a(t+2\Delta))F_{E}(j_{0})F_{E}(k_{0})\\ &+\sum_{i=0}^{i_{0}}\sum_{j=0}^{j_{0}}\sum_{k=0}^{k_{0}}f_{1}(i;a(t+2\Delta))e(j)e(k)\\ \times(1-&F_{2}(t-(i+j+k+n+2)\Delta;i+j+k+n,h))\end{split}$
@@
For completeness, we include the following converse result, which is a variation of (li2021close, , Theorem 3.6) due to the slightly different definition of safety violation here.
@@
Given a sufficiently large $s$, the private-mining attack violates the safety of a target transaction which becomes public during $(s,s+2\Delta]$ with at least the following probability:
(27) $\begin{split}&\left(1-\frac{a}{h}\right)e^{(h-a)(t+2\Delta)}\sum_{j=0}^{\infty}\sum_{k=0}^{\infty}e(j)\left(\frac{a}{h}\right)^{k}F_{1}(k;h(t+2\Delta))\\ &\;\times(1-F_{2}(t+\Delta-(j+k)\Delta;j+k+1,h))-e^{-h(t+2\Delta)}.\end{split}$
@@
5. Essential Elements of the Analysis
@@ 5.1. The Balancing Process
Two chains are said to agree on a target transaction if they both contain a common block which includes the transaction, or neither includes the transaction; otherwise, they are said to disagree on the target transaction. We also say blocks $b$ and $d$ agree (resp. disagree) on the target transaction if chains $b$ and $d$ agree (resp. disagree) on the target transaction.
Agreement about a target transaction is an equivalence relation, which typically divide all blocks into two equivalence classes. In Fig. 1, however, because transaction 4 is included in both blocks 2 and 3, its agreement divides all blocks into three equivalence classes, namely, $\{0,1,6\}$, $\{2,4,8,10\}$, and $\{3,5,7,9\}$.
@@
Given a target transaction, if an H-block (denoted as block $b$) is the first block mined on height $n$ and disagrees on the target transaction with any height-$(n-1)$ block mined before block $b$ becomes public, then block $b$ is said to be a balancing candidate, or simply a candidate. The period from the first time the disagreement arises to the time block $b$ becomes public is referred to as its candidacy. Evidently, if any disagreeing height-$(n-1)$ block is mined before block $b$, its candidacy begins from $t_{b}$.
@@
If during a balancing candidate’s candidacy, a disagreeing H-block is mined on the candidate’s height and no adversarial block is mined on that height, then the candidate is said to be balanced. Accordingly, its height is said to be a balanced height.
We use $X_{s,t}$ to denote the total number of balanced candidates whose candidacy is finished during $(s,t]$. With slight abuse notation, we let
(28) $X=\lim_{t\to\infty}X_{0,t}.$
@@
Figure 2. Two balancing candidates $b$ and $b^{\prime}$ at different times. Blocks on the same height are connected by dashed lines. Left: $b$ becomes a balancing candidate upon its mining; Right: $b^{\prime}$ becomes a candidate after $c$ is mined on a lower height to create a disagreement. The duration of each candidacy is marked red on the time axis. The duration of the first of those candidacies is exactly $\Delta$. The duration of the second one is strictly less than $\Delta$.
A candidacy is a small window of opportunity to balance the corresponding candidate (and thus create a balanced height). Figs. 2 illustrates two different balancing candidacies at different times. Examples of balancing can be found in Fig. 1. With regard to target transaction 4, blocks 4 and 5 balance each other on height 3; blocks 9 and 10 balance each other on height 5. No other heights are balanced. In particular, while H-blocks 3 and 6 are on the same height, they do not balance each other due to the existence of block 2 as an A-block on the same height. We shall show that the total number of heights that can be balanced (denoted as $X$) is essentially stochastically dominated by a geometric random variable.
5.2. Violation Event under Confirmation by Depth
@@
The pre-mining lead (of the adversary) at time $t$, denoted by $L_{t}$, is the height of the highest block mined by time $t$ minus the height of the highest H-block mined by time $t$.
The notion of pre-mining lead is to account for the potential lead the adversary may already have when the target transaction becomes public at around time $s$. In this work, $s$ is given so that the adversary cannot mount an adaptive attack (by influencing $s$) depending on the lead. Nonetheless, the adversary can try to create and maintain some lead (if possible), which is essentially characterized by a birth-death process.
@@
We say the genesis block is the 00-th jumper. After the genesis block, every jumper is the first H-block mined on its height. We use $J_{s,t}$ to denote the number of jumpers mined in $(s,t]$.
In the remainder of this section, we assume $s$ is a given time. We focus on the case of confirmation by depth $k$.
It is easy to see that the height growth of H-blocks is no less than the number of jumpers mined. Intuitively, for the transaction’s safety to be violated, there must exists an interval, during which the number of jumpers mined is fewer than the sum of the lead, the number of balanced heights, and the number of A-blocks mined. This is made precise in the following lemma.
Let $\tau$ denote the first time a block is mined that is $k$ or more blocks higher than the highest public block at time $s+3\Delta$.
@@
Suppose a target transaction becomes public during $(s,s+2\Delta]$. If this target transaction’s safety is violated, then the following event must occur:
(29) $L_{s}+X_{s,\tau}+\sup_{d\geq\tau}\{A_{s,d}-J_{\tau,d-\Delta}\}\geq k.$
@@
Suppose target transaction is public at time $s^{\prime}$, block $b^{\prime}$ is the highest public block at time $s^{\prime}+\Delta$. Let $\tau^{\prime}$ denote the first time a block is mined at height $h_{b^{\prime}}+k$ or above. The highest public H-block at time $d$ is at least $h_{b^{\prime}}+k+J_{\tau^{\prime},d-\Delta}$. If a violation occurs at time $d$, there must exist two chains disagreeing with each other on the target transaction, where both chains are no shorter than $h_{b^{\prime}}+k+J_{\tau^{\prime},d-\Delta}$. At time $s^{\prime}$, since the highest H-block is no higher than $h_{b^{\prime}}$, the highest A-block is no higher than $h_{b^{\prime}}+L_{s^{\prime}}$. Since $X_{s^{\prime},d}$ heights are balanced, A-blocks need to be mined at least on the remaining $k-X_{s^{\prime},d}$ heights after $h_{b^{\prime}}+L_{s^{\prime}}$. Along with A-blocks mined during $(s^{\prime},d]$, a violation implies the following,
(30) $h_{b^{\prime}}+L_{s^{\prime}}+X_{s^{\prime},d}+A_{s^{\prime},d}$ $\geq h_{b^{\prime}}+k+J_{\tau^{\prime},d-\Delta}.$
By Definition 18 of the pre-mining lead,
(31) $L_{s}+X_{s,d}+A_{s,d}\geq L_{s^{\prime}}+X_{s^{\prime},d}+A_{s^{\prime},d}.$
By definition of $\tau$, we have
(32) $k+J_{\tau^{\prime},d-\Delta}\geq k+J_{\tau,d-\Delta}.$
One way a violation occurs is for a balancing candidate to arise (regardless of whether the height is subsequently balanced). Thus, if a violation occurs at time $d>\tau^{\prime}$, we must have
(33) $X_{\tau^{\prime},d}=0.$
Thus, we have
(34) $L_{s}+X_{s,\tau}+A_{s,d}\geq k+J_{\tau,d-\Delta}.$
Therefore, a violation can only occur if (29) occurs. ∎
@@
5.3. A Sketch of the Statistical Analysis

### v5

@@ IV-B Throughput and Related Trade-offs
Empirically, the block propagation delay is approximately an affine function of the block size [21, 12] (even with a coded design [23]). That is, the propagation delay bound can be modeled as
$\Delta(B)=\frac{B}{r}+\nu$ (22)
where $B$ is the block size in KB and $r$ and $\nu$ are constants. Here, $r$ represents a fundamental limitation on the end-to-end transmission rate of the peer-to-peer network. Based on (21), this rate and the desired fault tolerance determine an upper bound for the optimistic throughput:
$\frac{\lambda B}{1+\lambda\Delta}<\frac{1-2\beta}{1-\beta-\beta^{2}}\,r.$ (23)
With a close upper bound on the probability of safety violation as a function of the system parameters and the confirmation latency, e.g., in the form of $p(\lambda\Delta,\beta,k)$, a node may select the confirmation depth $k$ to achieve a desired safety. In particular, the creator of a highly-valued transaction will most likely tolerate a longer-than-usual confirmation latency. The theorems in this paper allow this choice to be made much more accurately than before.
@@
$\text{maximize}_{\lambda,B,k}\quad$ $\frac{\lambda B}{1+\lambda\,\Delta(B)}$ (24a)
subject to $p(\lambda\,\Delta(B),\beta,k)\leq q$ (24b)
$k\leq\frac{(1-\beta)\lambda}{1+(1-\beta)\lambda\Delta}d$ (24c)
where $q$ denotes the maximum tolerable probability of violation and $d$ denotes the expected confirmation time. Here (24c) constrains $k$ such that at the slowest growth rate, gaining $k$ heights takes no more than $d$ units of time on average. In Sec. VI, we give such a numerical example under assumption (22) with empirically determined constants.
IV-C Bounds for Given Confirmation Time
Theorem 8 (safety guarantee for given confirmation time).
Suppose the mining rates and the delay upper bound satisfy (1). Given a sufficiently large $s$, the safety of a block mined by time of $s$ and is confirmed by time of $t$ can be violated with a probability no higher than
$\inf_{n\in\{1,2,\dots\}}\{1-F(n-1)+G(n,t)\},$ (25)
@@ IV-C Bounds for Given Confirmation Time
$G(n,t)=\sum_{i=0}^{\infty}\sum_{j=0}^{\infty}\sum_{k=0}^{\infty}f_{1}(i;a(t+2\Delta))e(j)e(k)\times(1-F_{2}(t-(i+j+k+n+2)\Delta;i+j+k+n,h))$ (26)
@@
$\begin{split}\overline{G}(n,t)&=1-F_{1}(i_{0};a(t+2\Delta))F_{E}(j_{0})F_{E}(k_{0})\\ &+\sum_{i=0}^{i_{0}}\sum_{j=0}^{j_{0}}\sum_{k=0}^{k_{0}}f_{1}(i;a(t+2\Delta))e(j)e(k)\times(1-F_{2}(t-(i+j+k+n+2)\Delta;i+j+k+n,h))\end{split}$ (27)
@@
For completeness, we include the following converse result, which is a variation of [6, Theorem 3.6] due to the slightly different definition of safety violation here.
@@
Given a sufficiently large $s$, the private-mining attack violates the safety of a target block mined by time $s$ and confirmed by time of $t$ with at least the following probability:
$\left(1-\frac{a}{h}\right)e^{(h-a)t}\sum_{j=0}^{\infty}\sum_{k=0}^{\infty}e(j)\left(\frac{a}{h}\right)^{k}F_{1}(k;ht)\times(1-F_{2}(t-(j+k)\Delta;j+k+1,h))-e^{-h(t-\Delta)}.$ (28)
@@
V Essential Elements of the Analysis
@@ V-A The Balancing Process
Two chains that are no lower than a target block are said to agree on the target block if either they both include the target block or neither includes it; otherwise, they are said to disagree on the target block.
Agreement on a target block is an equivalence relation, which divides all blocks higher than the target block into two equivalence classes. In Fig. 1, agreement about target block 3 divides blocks on height 2 and above into two equivalence classes, namely, blocks $\{3,5,7,9\}$ and blocks $\{2,4,6,8,10\}$.
@@
Given a target block, if an H-block (denoted as block $b$) is the first block mined on height $n$ and disagrees on the target block with any height-$(n-1)$ block mined before block $b$ becomes public, then block $b$ is said to be a balancing candidate, or simply a candidate. The period from the first time the disagreement arises to the time block $b$ becomes public is referred to as its candidacy. Evidently, if any disagreeing height-$(n-1)$ block is mined before block $b$, its candidacy begins from $t_{b}$.
@@
If a disagreeing H-block is mined on the candidate’s height and no A-block is mined on that height, then the candidate is said to be balanced. Accordingly, its height is said to be a balanced height.
Given a target block $g$, we use $X_{s,t}^{g}$ to denote the total number of balanced candidates created during $(s,t]$. We also define $X^{g}_{t}=X^{g}_{t_{g},t}$ and
$X^{g}=\lim_{t\to\infty}X_{t}^{g}.$ (29)
@@
Figure 2: Balancing candidates $b$ and $b^{\prime}$ for target block $g$ and $g^{\prime}$, respectively. Blocks on the same height are connected by dashed lines. Left: block $b$ becomes a balancing candidate upon its mining; its candidacy is $[t_{b},t_{b}+\Delta]$. Right: block $b^{\prime}$ becomes a candidate after a lower block $c$ is mined, in disagreement with block $b^{\prime}$. Block $b^{\prime}$ has a shorter candidacy $[t_{c},t_{b^{\prime}}+\Delta]$.
A candidacy is a small window of opportunity to balance the corresponding candidate (and thus create a balanced height). Figs. 2 illustrates two different balancing candidacies at different times. Examples of balancing can be found in Fig. 1. If block 3 is the target, blocks 4 and 5 balance each other on height 3; blocks 9 and 10 balance each other on height 5. No other heights are balanced. In particular, while H-blocks 3 and 6 are on the same height, they do not balance each other due to the existence of block 2 as an A-block on the same height. We shall show that the total number of heights that can be balanced is essentially stochastically dominated by a geometric random variable.
V-B Violation Event under Confirmation by Depth
@@
The pre-mining lead (of the adversary) at time $t$, denoted by $L_{t}$, is the height of the highest block mined by time $t$ minus the public height at time $t$.
The notion of pre-mining lead is to account for the potential lead the adversary may already have when the target block is included by a blockchain. In this work, $s$ is given so that the adversary cannot mount an adaptive attack (by influencing $s$) depending on the lead. Nonetheless, the adversary can try to create and maintain some lead (if possible), which is essentially characterized by a birth-death process.
@@
The first H-block mined on its height is called a jumper. We use $J_{s,t}$ to denote the number of jumpers mined in $(s,t]$.
In the remainder of this section, we assume $s$ is a given time. We focus on the case of confirmation by depth $k$. It is easy to see that the height growth of H-blocks is no less than the number of jumpers mined. Intuitively, for a block’s safety to be violated, there must exist an interval, during which the number of jumpers mined is fewer than the sum of the pre-mining lead, the number of balanced heights, and the number of A-blocks mined. This is made precise in the following lemma.
@@
Given $s$, let $\tau$ denote the arrival time of the first H-block that is at least $k$ higher than the public height at time $s$. If a block $g$ extends an $s$-credible chain and its safety is violated, then the following must hold:
$L_{s}+X_{t_{f}+\Delta}^{g}+\sup_{d\geq\tau}\{A_{s,d}-J_{\tau,d-\Delta}\}\geq k$ (30)
where block $f$ is the first block mined on height $h_{g}+k-1$, whose mining time is denoted as $t_{f}$.
@@
We first show that (30) is implied by a slightly stronger inequality:
$L_{s}+X_{t_{f}+\Delta}^{g}+\sup_{d\geq t_{f}}\{A_{s,d}-J_{\tau,d-\Delta}\}\geq k,$ (31)
where the domain of optimization $d\geq\tau$ in (30) is replaced by $d\geq t_{f}$. In the case of $t_{f}\geq\tau$, the domain in (30) is larger, yielding a higher supremum value, hence (30) holds as long as (31) holds. In the remaining case of $t_{f}<\tau$, we have
$\sup_{d\geq t_{f}}\{A_{s,d}-J_{\tau,d-\Delta}\}$ $=\max\left\{\sup_{d>\tau}\{A_{s,d}-J_{\tau,d-\Delta}\},\sup_{\tau\geq d\geq t_{f}}\{A_{s,d}-J_{\tau,d-\Delta}\}\right\}$ (32)
$=\max\left\{\sup_{d>\tau}\{A_{s,d}-J_{\tau,d-\Delta}\},A_{s,\tau}\right\}$ (33)
$=\sup_{d\geq\tau}\{A_{s,d}-J_{\tau,d-\Delta}\}$ (34)
where in (33) and (34) we have both used the fact that $J_{\tau,d-\Delta}=0$ for $d\leq\tau$. Hence, in this case (30) and (31) are equivalent.
The remainder of this proof is devoted to establishing (31). Here we let all agreement relationship be with respect to block $g$. Evidently, block $g$ can not be $k$-confirmed before $t_{f}$. Let chain $b$ denote a highest public chain by time $s$. Since block $g$ extends an $s$-credible chain, we have $h_{g}>h_{b}$ and hence
$h_{f}\geq h_{b}+k$ (35)
by definition of block $f$ in the lemma. Let block $c$ denote a highest H-block mined by $t_{f}$. Let chain $z$ denote a highest chain that disagrees with chain $c$ at time $t_{f}$. Definitions of some frequently used notations in this proof are listed in Table II.
Notation Meaning
$s$ a given time for commitment rule
chain $b$ a highest public chain by $s$
block $c$ a highest H-block mined by $t_{f}$
block $f$ the first block mined on height $h_{g}+k-1$
block $g$ the target block
chain $z$ a highest chain that disagrees with chain $c$ at $t_{f}$
$\tau$ the arrival time of the first H-block that is at least $k$ higher than chain $b$
TABLE II: Frequently used notations in the proof of Lemma 10.
We first establish the following upper bound on the number of jumpers mined during $(\tau,t_{f}]$:
$J_{\tau,t_{f}}\leq h_{f}-(h_{b}+k).$ (36)
If $\tau\geq t_{f}$, $J_{\tau,t_{f}}=0$ by convention, so that (36) holds trivially due to (35). If $\tau<t_{f}$, the jumpers between $\tau$ and $t_{f}$ must be between heights $h_{b}+k+1$ and $h_{f}$, hence (36) must hold.
We establish (31) in two separate cases depending on whether $h_{c}\leq h_{b}+L_{s}$.
  • Case 1)
    $h_{c}\leq h_{b}+L_{s}$. In this case no H-blocks higher than $h_{b}+L_{s}$ are mined by $t_{f}$, so there is at least one A-block on each height from $h_{b}+L_{s}+1$ to $h_{f}$ mined between $s$ and $t_{f}$, i.e.,
    $A_{s,t_{f}}$ $\geq h_{f}-(h_{b}+L_{s})$ (37)
    $\geq k+J_{\tau,t_{f}}-L_{s},$ (38)
    where (38) is due to (36). Since $X_{t_{f}+\Delta}^{g}\geq 0$ and $\sup_{d\geq t_{f}}\{A_{s,d}-J_{\tau,d-\Delta}\}\geq A_{s,t_{f}}-J_{\tau,t_{f}-\Delta}\geq A_{s,t_{f}}-J_{\tau,t_{f}}$, (38) implies (31).
  • Case 2)
    $h_{c}>h_{b}+L_{s}$. We divide the proof in this case into two parts (2-a) and (2-b), where the first part focuses on blocks mined during $(s,t_{f}]$, and the second part focuses on blocks mined after $t_{f}$.
    • Part 2-a)
      Since chains $c$ and $z$ disagree, there exist at least two disagreeing chains of height $\min\{h_{c},h_{z}\}$ by $t_{f}$. This implies that at least one A-block is mined during $(s,t_{f}]$ on every height between $h_{b}+L_{s}+1$ and $\min\{h_{c},h_{z}\}$ except those heights that are balanced by time $t_{f}$. The number of balanced heights is no larger than $X_{t_{f}}^{g}$. In addition, since no H-block is higher than $h_{c}$, at least one A-block is mined on each height from $h_{c}+1$ to $h_{f}$. Therefore, the number of A-blocks is lower bounded:
      $A_{s,t_{f}}\geq(\min\{h_{c},h_{z}\}-(h_{b}+L_{s})-X_{t_{f}}^{g})+(h_{f}-h_{c}).$ (39)
    • Part 2-b)
      We now focus on blocks mined after $t_{f}$ in two different cases depending on the relative positions of blocks $c$ and $z$.
      • Case $h_{z}\geq h_{c}$.
        Since $\min\{h_{c},h_{z}\}=h_{c}$, (39) reduces to
        $A_{s,t_{f}}$ $\geq h_{f}-(h_{b}+L_{s})-X_{t_{f}}^{g}.$ (40)
        Comparing (36) and (40) yields
        $L_{s}+X_{t_{f}}^{g}+A_{s,t_{f}}-J_{\tau,t_{f}}$ $\geq k.$ (41)
        Noting that $X^{g}_{t_{f}}\leq X^{g}_{t_{f}+\Delta}$, $J_{\tau,t_{f}}\geq J_{\tau,t_{f}-\Delta}$, and reducing the domain $d\geq t_{f}$ in (31) to $d=t_{f}$ reduces the supremum, we have established (31) based on (41) for Case $h_{z}\geq h_{c}$.
      • Case $h_{z}<h_{c}$.
        In this case, block $c$ must agree with block $f$; otherwise chain $z$ cannot be the highest to disagree with chain $c$. Since block $f$ is the first block that is $k-1$ higher than block $g$, whose safety is subsequently violated, the first chain that is credible and also disagrees with chains $c$ and $f$ must be created at some time $d\geq t_{f}$ to create a safety violation. All jumpers mined during $(t_{f},d-\Delta]$ become public by time $d$, so the highest public block at time $d$ is at least as high as $h_{c}+J_{t_{f},d-\Delta}$. Because the highest disagreeing chain at $t_{f}$ is no higher than $h_{z}$, those blocks of the $d$-credible disagreeing chain between heights $h_{z}+1$ and $h_{c}+J_{t_{f},d-\Delta}$ must be mined during $(t_{f},d]$. Moreover, after $t_{f}+\Delta$, a violation occurs as soon as a balancing candidate is created (before it is even balanced). Hence at most $X_{t_{f},t_{f}+\Delta}^{g}$ of heights $(h_{z}+1,\dots,h_{c}+J_{t_{f},d-\Delta})$ can be balanced, whereas at least one A-block is mined on the remaining heights. Putting it altogether, we have
        $A_{t_{f},d}$ $\geq h_{c}+J_{t_{f},d-\Delta}-h_{z}-X_{t_{f},t_{f}+\Delta}^{g}.$ (42)
        Adding (39) with $\min\{h_{c},h_{z}\}=h_{z}$ and canceling $h_{z}$ and $h_{c}$ on both sides, we have
        $A_{s,d}$ $\geq J_{t_{f},d-\Delta}+h_{f}-(h_{b}+L_{s})-X_{t_{f}+\Delta}^{g}$ (43)
        $\geq J_{t_{f},d-\Delta}+J_{\tau,t_{f}}+k-L_{s}-X_{t_{f}+\Delta}^{g}$ (44)
        $=J_{\tau,d-\Delta}+k-L_{s}-X_{t_{f}+\Delta}^{g},$ (45)
        where (44) is due to (36). Since $\sup_{d\geq t_{f}}\{A_{s,d}-J_{\tau,d-\Delta}\}$ in (31) is greater or equal to $A_{s,d}-J_{\tau,d-\Delta}$, (45) implies (31).
We caution that Lemma 10 is very general and at the same time quite delicate. While block $b$ in the proof is the highest public block by $s$, block $g$ can be arbitrarily higher than height $h_{b}+1$, and may or may not be mined after $s$. In fact, even block $f$ may or may not be mined before $s$, as there is the possibility that chain $f$ already $k$-confirms block $g$, yet all of those confirming blocks including block $g$ itself are nonpublic. We illustrate (30) in Lemma 10 using a pair of concrete chains in Fig 3.
$d$$t_{f}$$\tau$$s$$b$$g$$c$$e$$z$$v$
Figure 3: An example of (30) in Lemma 10. Let the depth of confirmation $k=5$. Blocks $b$ and $e$ are on the same height. Block $b$ is the highest public chain at $s$ (block $g$’s parent is not yet public). The lead $L_{s}=2$. Here $c$ and $f$ refer to the same block, i.e., $c=f$. At time $t_{f}$, block $g$ is $k$-confirmed by the highest chain $c$, as $h_{c}=h_{g}+k-1$. The highest disagreeing chain $z$ at $t_{f}$ is on height $h_{z}=h_{g}+2<h_{c}$. The (first) violation occurs at time $d=t_{v}$, when chain $v$ is mined, which is credible as $h_{v}=h_{c}$. Height $h_{g}+1$ is balanced, with $X^{g}_{t_{f}}=1$. A single jumper is mined during $(\tau,d-\Delta]$. Three A-blocks are mined during $(s,d]$. Therefore, we have $L_{s}+X_{t_{f}+\Delta}^{g}+A_{s,d}-J_{\tau,d-\Delta}=5=k$, i.e., (30) holds in this case.
@@
V-C A Sketch of the Statistical Analysis